OpenStudy (jazzyfa30):

Write an equation for a circle with center at (2,3) and radius of 6.

OpenStudy (jazzyfa30):

@Jonask

OpenStudy (anonymous):

we need to write as \[(x-a)^2+(y-b)^2=r^2\] where (a,b) is center and r=6

OpenStudy (anonymous):

can you do that

OpenStudy (jazzyfa30):

ok so \[(2-a)^2+(3-b)^2=6\] like that

OpenStudy (anonymous):

substitue in the place of a and b but leave x and y

OpenStudy (jazzyfa30):

ummmm so \[(x-2)^2+(y-3)^2=6\] this way????

OpenStudy (anonymous):

yes but \[6^2\] on the right

OpenStudy (jazzyfa30):

oh ok

OpenStudy (anonymous):

yes no p

OpenStudy (jazzyfa30):

???????

OpenStudy (anonymous):

no problem

OpenStudy (jazzyfa30):

Write an equation for a circle with center at (-2,7) and diameter of 14.

OpenStudy (anonymous):

now we first find the radius ie half of the diameter 14/2=7=r then do the same procedure as before

OpenStudy (jazzyfa30):

ok so the answer is \[(x+2)^2+(y-7)^2=14^2\]

OpenStudy (anonymous):

yes but be care ful i explained above wat u shud do to 14

OpenStudy (jazzyfa30):

oh its 7^2

OpenStudy (anonymous):

yes good

OpenStudy (anonymous):

great job

OpenStudy (jazzyfa30):

this next one has a graph

OpenStudy (anonymous):

whe we count the blocks from center we see that the radius is r=3 center(3,-1)

OpenStudy (jazzyfa30):

ok the answer is \[(x-3)^2+(y+1)^2=3^2\]

OpenStudy (anonymous):

yes good

OpenStudy (jazzyfa30):

ok 1 more Use the equation below to find the center and radius of the circle. x2 -2x +y2 -2y -23 =0 Select one: a. (-1,-1) r=23 b. (1,-1) r=25 c. (1,1) r =5 d. (1,1) r= 23

OpenStudy (anonymous):

\[ x^2 -2x +y^2 -2y -23 =0\] you need to complete square

OpenStudy (anonymous):

\[(x -1)^2+(y-1)^2=23+1+1=25 \]

OpenStudy (anonymous):

center(1,1) r=5

OpenStudy (jazzyfa30):

ok thank u